.) The following data refer to quantity of sweets demanded by students in given school.
Price (Kshs) Quantity demanded
1 100

1.10 80

Required:
i) Calculate the price elasticity of demand considering a price increase form Ksh. 1.00 to Ksh. 1.10.
(3.5 marks)
ii) Calculate the price elasticity of demand when price decreases from Ksh. 1.10 to Ksh. 1.00. (3.5 marks)
1
Expert's answer
2019-10-29T09:56:43-0400

Solution:

Calculate the price elasticity of demand considering a price increase form Ksh. 1.00 to Ksh. 1.10.

Q2=80, Q1=100, p2=1.10, p1=1.00


"E= (Q2-Q1)/(p2-p1)\times (p2+p1)/(Q2+Q1)"

"E= (80-100)/(1.10-1.00) \times(1.10+1.00)/(80+100)≈-2.33"


Answer: -2.33

Calculate the price elasticity of demand when price decreases from Ksh. 1.10 to Ksh. 1.00.

Q2=100, Q1=80, p2=1.00, p1=1.10


"E= (100-80)/(1.00-1.10)\times (1.00+1.10)/(80+100)≈-2,33"

Answer: -2.33